OpenStudy (anonymous):

Find all solutions to the equation. sin2x + sin x = 0

OpenStudy (anonymous):

2sin(x)cos(x)+sin(x)=0 sin(x)[2cos(x)+1]=0 either sin(x)=0 x=n*pi or cos(x)=-1/2 x=2n*pi+ pi/6 , 2n*pi -pi/6

OpenStudy (anonymous):

Or, if you mean \[\sin^2x+\sin x=0\\ \sin x\left(\sin x+1\right)=0\] Then you have \[\begin{matrix}\sin x=0&&&\sin x=-1\\ x=2n\pi&&&x=\frac{3\pi}{2}+2n\pi\end{matrix}\]